# Analysis of mean field games via Fokker-Planck-Kolmogorov equations:

### v1

@@ 3. Auxiliary results
on the space $U\times\mathbb{R}^{d}\times[0,T]$. For every $t\in[0,T]$ the measure $\delta_{u(x,t)}(du)\mu_{t}(dx)$ is well defined since the function $x\mapsto\delta_{u(x,t)}(E)=I_{E}(u(x,t))$ is Borel measurable for every Borel set $E\subset U$. Let us proof that the family of measures $\bigl{(}\delta_{u(x,t)}(du)\mu_{t}(dx)\bigr{)}_{t\in[0,T]}$ is Borel measurable in $t$. Note that for every Borel set $C\subset U\times\mathbb{R}^{d}$ we have the equality
@@
where $I_{C}$ is the indicator of $C$. According to [11, Theorem 10.7.2], the mapping
@@
is Borel measurable. Thus the measure $\Pi$ is well defined. Conditional measures are discussed in [11, Chapter 10].
@@
Let $\sigma\in\mathcal{M}_{R,M}(V)$. Assume that $\nu$ is a Borel probability measure on $\mathbb{R}^{d}$ such that $V\in L^{1}(\nu)$. Denote by $S_{R}(\sigma)$ the set of measures $\Pi\in P_{R}$ such that the projection of $P$ on $(x,t)$ is a measure $\mu=\mu_{t}\,dt\in\mathcal{M}_{R}(V)$ satisfying the conditions: 1) $\mu_{0}=\nu$ and 2) the equality
@@
(ii) the projection of $\Pi\in S_{R}(\sigma)$ on $(x,t)$ belongs to the set $\mathcal{M}_{R,M}^{\omega}(V)$ for some $\omega=\{\omega_{\psi}\}$,
@@
$L_{\sigma,u_{0}}V(x)\leq C_{L}V(x)+h(|u_{0}|)+C_{L}Re^{Mt}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}.$
According to Remark 3.4, there exists a continuous mapping $t\mapsto\mu_{t}$ such that the measure $\mu=\mu_{t}\,dt$ is a solution to the Cauchy problem
@@
Moreover, one has the inequality
$\int_{\mathbb{R}^{d}}V(x)\,\mu_{t}(dx)\leq e^{C_{L}T}\Bigl{(}\int_{\mathbb{R}^{d}}V(x)\,d\nu+h(|u_{0}|)T+C_{L}T\beta_{V,W}(Re^{MT})Re^{MT}\Bigr{)}+\frac{1}{4}Re^{Mt},$
where we use the equality $M=5C_{L}$. There exists a number $R_{0}>0$ such that for all $R>R_{0}$ the estimates
$h(|u_{0}|)T\leq\gamma R,\quad e^{C_{L}T}\Bigl{(}\int_{\mathbb{R}^{d}}V(x)\,d\nu+h(|u_{0}|)T+C_{L}T\beta_{V,W}(Re^{MT})Re^{MT}\Bigr{)}\leq\frac{3R}{4}$
@@
Let the support of $\psi$ be in some open ball $B$. According to the condition (H1.1), the functions $a^{ij}$, $b^{i}$ and $q^{im}$ are bounded on $B\times[0,T]$. It follows that for a number $C>0$ depending on the function $\psi$, the ball $B$ and the coefficients $a^{ij}$, $b^{i}$ and $q^{im}$ we have the estimate
@@
Therefor $\mu\in\mathcal{M}_{R}^{\omega}(V)$, where $\omega=\{\omega_{\psi}\}$.
@@
According to the condition (H1.3), the equality $\lim_{n\to\infty}C_{n}(t)=0$ holds for all $t\in[0,T]$. By the condition (H1.1) we have $\sup_{n,t}C_{n}(t)<\infty$. Moreover, the exists a number $\widetilde{C}>0$ such that $|u|\leq\widetilde{C}+\widetilde{C}h(|u|)$ for all $u\in U$. Hence we derive the estimate
$\int_{U\times\mathbb{R}^{d}\times[0,t]}\Bigl{|}L_{\sigma^{n},u}\psi-L_{\sigma,u}\psi\Bigr{|}\,d\Pi^{n}\leq\widetilde{C}\Bigl{(}1+\gamma R\Bigr{)}\sup_{x}\Bigl{(}|\nabla\psi(x)|+\|D^{2}\psi(x)\|\Bigr{)}\int_{0}^{t}C_{n}(s)\,ds.$
Since the sequence
$\widetilde{C}\Bigl{(}1+\gamma R\Bigr{)}\sup_{x}\Bigl{(}|\nabla\psi(x)|+\|D^{2}\psi(x)\|\Bigr{)}\int_{0}^{t}C_{n}(s)\,ds$
tends to zero, it suffices to verify the equality
@@
Finally, we need to pass to the limit in the integral of the function $\langle Q(x,t,\sigma)u,\nabla\psi(x)\rangle$ with respect to $\mu_{t}^{n}\,dt$. Let $\zeta_{N}\in C(\mathbb{R}^{d_{1}})$, $0\leq\zeta_{N}\leq 1$, $\zeta_{N}(u)=1$ if $|u|<N$ and $\zeta_{N}(u)=0$ if $|u|>2N$. Set $u^{N}=u\zeta_{N}(u)$. Using the inequality $|u-u^{N}|\leq 2|u|I_{|u|>N}$, we obtain the estimate
@@
As in Remark 3.4 we cnsider the Cauchy problem
@@
Then, according to [14, Theorem 9.4.3], the class of sub–probability solutions contains precisely one element $\mu=\mu_{t}\,dt$.
In the following lemma we prove that for a given sequence $\sigma^{n}$ converging to $\sigma$ and some measure $\Pi\in S_{R}(\sigma)$ there exists a sequence of measures $\Pi^{n}\in S_{R}(\sigma^{n})$ such that the measures $\Pi^{n}$ converge weakly to the measure $\Pi$. This assertion plays a crucial role in the proof of Theorem 2.1. The main difficulty is that we need to solve degenerate Fokker–Planck–Kolmogorov equations with irregular coefficients. A similar problem arises in the proof of the Ambrosio–Figalli–Trevisan superposition principle [15]. Therefor the part of the proof repeats the arguments from [15]. The method of doubling variables is used. This method seems new for the Fokker–Planck–Kolmogorov equations but it is just an analytical form of a known method in stochastic differential equations (see [34, Chapter 4]).
@@
There exists a number $R_{0}$ such that for every $R>R_{0}$ the following statement is fulfilled. Suppose that $\sigma^{n},\sigma\in\mathcal{M}_{R,M}^{\omega}(V)$ and the measures $\sigma^{n}$ converge VV–weakly to the measure $\sigma$. Then for every measure $\Pi\in S_{R}(\sigma)$ of the form
@@
where $(x,t)\mapsto u(x,t)$ is a Borel function, there exists a sequence of measures $\Pi^{n}\in S_{R}(\sigma^{n})$ such that for every $n$ the projection of $\Pi^{n}$ on $(u,t)$ is equal to the projection of $\Pi$ on $(u,t)$ and the measures $\Pi^{n}$ converge weakly to the measure $\Pi$.
@@
$\|A(y,t,\sigma)\|+|b(y,t,\sigma)|+h^{*}(\|Q(y,x,\sigma)\|)\leq C_{\sigma}V(y).$
@@
Let us verify that measures $\Pi^{n}$ converge weakly to the measure $\Pi$. Since $\Pi^{n},\Pi\in P_{R}$ and $P_{R}$ is a compact set, it suffices to prove that if the measures $\Pi^{n}$ converge, then the limit is equal to $\Pi$. To do this it suffices to extract a subsequence converging to the measure $\Pi$.
@@
$\Bigl{|}\int_{0}^{T}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle Q(x,t,\sigma)(u(y,t)-\widetilde{u}(y,t)),\nabla_{x}\psi(x,y)\rangle(\pi_{t}^{n}+\pi)(dxdy)\,dt\Bigr{|}\leq\\ \leq 2C(\psi)\sup_{x\in B,t\in[0,T]}\|Q(x,t)\|\int_{0}^{T}\int_{\mathbb{R}^{d}}|u(y,t)-\widetilde{u}(y,t)|\mu_{t}(dy)\,dt.$
@@
Letting $\delta\to 0+$, we conclude that the equality $x=y$ holds for $\pi_{t}$ — almost all $(x,y)$. The idea of this proof, in particular, the choice of $H$, is well known (see, for example, [43]).
@@
Let $0<\varepsilon<\min\{2^{-1},T-T^{\prime}\}$ and $\phi$ be the standard Gaussian density on $\mathbb{R}^{d}$. We also use the function $\omega_{\varepsilon}(y)=\varepsilon^{-d}\omega(y/\varepsilon)$, where $\omega\in C_{0}^{\infty}(\mathbb{R}^{d})$, $\omega\geq 0$, $\|\omega\|_{L^{1}}=1$ and the support of $\omega$ is in $\{y\colon\,|y|<1\}$. Set
$\mu_{t}^{\varepsilon}(y)=\varepsilon\phi(y)+\frac{1-\varepsilon}{\varepsilon}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$a^{ij}_{\varepsilon}(y,t)=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}a^{ij}(z,s,\sigma)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$b^{i}_{\varepsilon}(y,t)=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}b^{i}(z,s,\sigma)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$(Qu)^{i}_{\varepsilon}(y,t)=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}Q(z,s,\sigma)u(s,z)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$u_{\varepsilon}(t,y)=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}u(s,z)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
@@
where the operator $L_{\sigma,u(z,s)}$ acts on the function $z\mapsto\omega_{\varepsilon}(y-z)$. Note that
@@
are integrable on $\mathbb{R}^{d}\times[0,T^{\prime}]$ with respect to the measure $\mu_{t}^{\varepsilon}(y)\,dy\,dt$ and $L^{1}$–norms of these functions are majorized by a constant independent on $\varepsilon$. Let us prove this observation. Consider the function $(Qu)^{i}_{\varepsilon}$. We have
@@
$\mathbb{B}_{n}^{\varepsilon,j}(x,y,t)=\left(\begin{array}[]{l}\zeta_{j}(x)b(x,t,\sigma^{n})+\zeta_{j}(x)Q(x,t,\sigma^{n})u_{\varepsilon}(y,t)\\ b_{\varepsilon}(y,t,\sigma)+(Qu)_{\varepsilon}(y,t,\sigma)-y\frac{\varepsilon\phi(y)}{\mu_{t}^{\varepsilon}(y)}\\ \end{array}\right).$
@@
over all probability measures $\eta$ with projections $\nu$ and $\nu^{\varepsilon}$ on the factors. It is known (see, for instance [13]) that the minimum of this functional tends to zero as $\varepsilon\to 0$ because the measures $\pi_{0}^{\varepsilon}$ converge weakly to the measure $\nu\circ(x,x)^{-1}$.
There exists a measure $\pi^{n,\varepsilon,j}=\pi_{t}^{n,\varepsilon,j}\,dt$ such that this measure is given by a family of sub–probability measures $\pi_{t}^{n,\varepsilon,j}$ and $\pi^{n,\varepsilon,j}$ is a solution to the Cauchy problem with the operator $\mathbb{L}_{n}^{\varepsilon,j}$ and the initial condition $\pi_{0}^{\varepsilon}$.
@@
Let us consider the function $\zeta_{N}$ as above and arbitrary function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$. Substitute the function $\zeta_{N}(x)\psi(y)$ into the integral identity which defines the solution $\pi^{n,\varepsilon,j}$. Letting $N\to\infty$, we obtain that the projection of the measure $\pi^{n,\varepsilon,j}$ on $y$ is a solution to the Cauchy problem for the Fokker–Planck–Kolmogorov equation with the operator $L_{\varepsilon}$ and the initial condition $\nu^{\varepsilon}$. Since the coefficients are integrable with respect to the measure $\mu^{\varepsilon}_{t}(y)\,dy$ and the matrix $A_{\varepsilon}$ is locally Lipschitzian in $y$ and locally non-degenerate, the class of sub–probability solutions contains precisely one element $\mu^{\varepsilon}=\mu_{t}^{\varepsilon}(y)\,dy\,dt$. It follows that the projection of $\pi_{t}^{n,\varepsilon,j}$ on $y$ is equal to the measure $\mu_{t}^{\varepsilon}(y)\,dy$. In particular, $\pi_{t}^{n,\varepsilon,j}$ is a probability measure.
For every $t\in[0,T^{\prime}]$ let denote by $\mu^{n,\varepsilon,j}_{t}$ the projection of the measure $\pi_{t}^{n,\varepsilon,j}$ on $x$. Substitute the function $V(x)\zeta_{N}(x)\zeta_{N}(y)$ into the integral identity which defines the solution $\pi^{n,\varepsilon,j}$. Note that the coefficients depending on $x$ vanish outside a ball of sufficiently large radius. Moreover, the coefficients depending on $y$ are integrable with respect to the measure $\pi^{n,\varepsilon,j}$ since the projection of $\pi^{n,\varepsilon,j}$ on $(y,t)$ equals the measure $\mu_{t}^{\varepsilon}(y)\,dydt$). Letting $N\to\infty$, we obtain the inequality
@@
$\zeta_{j}(x)L_{\sigma^{n},u(y,s)}V(x,s)\leq C_{L}V(x)+C_{L}Re^{Ms}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}+h(|u_{\varepsilon}(y,s)|).$
We need the following version of Jensen’s inequality. Assume that $\Phi$ is a convex nonnegative function on $0,+\infty)$ and $\Phi(0)=0$. Let $\xi\geq 0$ be a measurable function on a measurable space $(X,\mathcal{X})$ with a sub–probability measure $\omega$. Then
@@
Let us proof this estimate. If $\omega(X)=0$, then the inequality trivially holds since $\Phi(0)=0$. Let $\omega(X)>0$. The convexity of $\Phi$ and the equality $\Phi(0)=0$ imply the inequality $\Phi(\lambda v)\leq\lambda\Phi(v)$ for every $v\geq 0$ and $0<\lambda<1$. Thus we obtain
@@
Taking into account the last estimate and the fact that the projection of $\pi^{n,\varepsilon,j}_{t}$ on $y$ does not depend on $j$, we obtain that for every $t$ the sequence $\pi^{n,\varepsilon,j}_{t}$ is tight. Hence for every $t$ one can extract a convergent subsequence $\pi^{n,\varepsilon,j_{m}}_{t}$. Arguing as in Lemma 3.3, we extract a subsequence $j_{m}$ such that for every $t$ the measures $\pi^{n,\varepsilon,j_{m}}_{t}$ converge weakly to a probability measure $\pi^{n,\varepsilon}_{t}$. Moreover, the mapping $t\mapsto\pi^{n,\varepsilon}_{t}$ is continuous with respect to the weak topology. For every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ there exists a number $m_{0}$ such that for all $m>m_{0}$ and for every $(x,y,t)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\times[0,T^{\prime}]$ one has the equality

### v2

@@ 3. Auxiliary results
on the space $U\times\mathbb{R}^{d}\times[0,T]$. For every $t\in[0,T]$ the measure $\delta_{u(x,t)}(du)\mu_{t}(dx)$ is well defined since the function $x\mapsto\delta_{u(x,t)}(E)=I_{E}(u(x,t))$ is Borel measurable for every Borel set $E\subset U$. Let us prove that the family of measures $\bigl(\delta_{u(x,t)}(du)\mu_{t}(dx)\bigr)_{t\in[0,T]}$ is Borel measurable in $t$. Note that for every Borel set $C\subset U\times\mathbb{R}^{d}$ we have the equality
@@
where $I_{C}$ is the indicator of $C$. According to [12, Theorem 10.7.2], the mapping
@@
is Borel measurable. Thus the measure $\Pi$ is well defined. Conditional measures are discussed in [12, Chapter 10].
@@
Let $\sigma\in\mathcal{M}_{R,M}(V)$. Assume that $\nu$ is a Borel probability measure on $\mathbb{R}^{d}$ such that $V\in L^{1}(\nu)$. Denote by $S_{R}(\sigma)$ the set of measures $\Pi\in P_{R}$ such that the projection of $\Pi$ on $(x,t)$ is a measure $\mu=\mu_{t}\,dt\in\mathcal{M}_{R,M}(V)$ satisfying the conditions: 1) $\mu_{0}=\nu$ and 2) the equality
@@
(ii) there exists $\omega=\{\omega_{\psi}\}$ such that for every $\Pi\in S_{R}(\sigma)$ the projection of $\Pi$ on $(x,t)$ belongs to the set $\mathcal{M}_{R,M}^{\omega}(V)$,
@@
$L_{\sigma,u_{0}}V(x)\leq C_{L}V(x)+h(|u_{0}|)+C_{L}Re^{Mt}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT},$
where $\beta_{V,W}$ is defined in Remark 3.2. According to Remark 3.4, there exists a continuous mapping $t\mapsto\mu_{t}$ such that the measure $\mu=\mu_{t}\,dt$ is a solution to the Cauchy problem
@@
Applying the estimates from Remark 3.4 with $C=C_{L}$, $\mathcal{V}=V$ and
$\mathcal{W}(x,t)=h(|u_{0}|)+C_{L}Re^{Mt}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT},$
we obtain the inequality
$\int_{\mathbb{R}^{d}}V(x)\,\mu_{t}(dx)\leq e^{C_{L}T}\Bigl(\|V\|_{L^{1}(\nu)}+h(|u_{0}|)T+C_{L}T\beta_{V,W}(Re^{MT})Re^{MT}\Bigr)+\frac{1}{4}Re^{Mt}.$
Here we also use the equality $M=5C_{L}$. There exists a number $R_{0}>0$ such that for all $R>R_{0}$ the estimates
$h(|u_{0}|)T\leq\gamma R,\quad e^{C_{L}T}\Bigl(\|V\|_{L^{1}(\nu)}+h(|u_{0}|)T+C_{L}T\beta_{V,W}(Re^{MT})Re^{MT}\Bigr)\leq\frac{3R}{4}$
@@
Let the support of $\psi$ be in some open ball $B$. According to the condition (H1.1) we have
$C(B)=\sup_{x\in B,t\in[0,T],\theta\in\mathcal{M}_{R}(V)}\Bigl(\|A(x,t,\mu)\|+|b(x,t,\theta)|+\|Q(x,t,\theta)\|\Bigr)<\infty.$
Set
$C=C(B)\sup_{x}\Bigl(\|D^{2}\psi(x)\|+|\nabla\psi(x)|\Bigr).$
We have the estimate
@@
Therefor $\mu\in\mathcal{M}_{R,M}^{\omega}(V)$, where $\omega=\{\omega_{\psi}\}$.
@@
According to the condition (H1.3), the equality $\lim_{n\to\infty}C_{n}(t)=0$ holds for all $t\in[0,T]$. By the condition (H1.1) we have $\widetilde{C}=\sup_{n,t}C_{n}(t)<\infty$. Hence we derive the estimate
$\int_{U\times\mathbb{R}^{d}\times[0,t]}\Bigl|L_{\sigma^{n},u}\psi-L_{\sigma,u}\psi\Bigr|\,d\Pi^{n}\leq\sup_{x}\Bigl(|\nabla\psi(x)|+\|D^{2}\psi(x)\|\Bigr)\int_{U\times\mathbb{R}^{d}\times[0,t]}C_{n}(s)(1+|u|)\,d\Pi^{n}.$
Since for every number $N\geq 1$ the estimate
$\int_{U\times\mathbb{R}^{d}\times[0,t]}C_{n}(s)(1+|u|)\,d\Pi^{n}\leq(1+N)\int_{0}^{t}C_{n}(t)\,dt+\gamma R\widetilde{C}\sup_{v>N}\frac{v}{h(v)}.$
is fulfilled and $\lim_{N\to\infty}\sup_{v>N}\frac{v}{h(v)}=0$, we obtain
$\lim_{n\to\infty}\int_{U\times\mathbb{R}^{d}\times[0,t]}\Bigl|L_{\sigma^{n},u}\psi-L_{\sigma,u}\psi\Bigr|\,d\Pi^{n}=0.$
Thus it suffices to verify the equality
@@
Finally, we need to pass to the limit in the integral of the function $\langle Q(x,t,\sigma)u,\nabla\psi(x)\rangle$ with respect to $\Pi^{n}$. Let $\zeta_{N}\in C(\mathbb{R}^{d_{1}})$, $0\leq\zeta_{N}\leq 1$, $\zeta_{N}(u)=1$ if $|u|<N$ and $\zeta_{N}(u)=0$ if $|u|>2N$. Set $u^{N}=u\zeta_{N}(u)$. Using the inequality $|u-u^{N}|\leq 2|u|I_{|u|>N}$, we obtain the estimate
@@
Note that the function $\omega_{\psi}(v)=Cv+Cvh^{-1}(\gamma R/v)$ from the statement (ii) of Lemma 3.2 does not depend on $\sigma$.
@@
As in Remark 3.4 we consider the Cauchy problem
@@
Then, according to [15, Theorem 9.4.3], the class of sub–probability solutions contains precisely one element $\mu=\mu_{t}\,dt$.
In the following lemma we prove that for a given sequence $\sigma^{n}$ converging to $\sigma$ and some measure $\Pi\in S_{R}(\sigma)$ there exists a sequence of measures $\Pi^{n}\in S_{R}(\sigma^{n})$ such that the measures $\Pi^{n}$ converge weakly to the measure $\Pi$. This assertion plays a crucial role in the proof of Theorem 2.1. The main difficulty is that we need to solve degenerate Fokker–Planck–Kolmogorov equations with irregular coefficients. A similar problem arises in the proof of the Ambrosio–Figalli–Trevisan superposition principle [16]. Therefor the part of the proof repeats the arguments from [16]. We use the method of doubling variables described in [9, Section 3.3]. Note that this method is just an analytical form of a known method in stochastic differential equations (see [46, Lemma 8.1.3], [36, Chapter 4] or [28]).
@@
There exists a number $R_{0}$ such that for every $R>R_{0}$ the following statement is fulfilled. Suppose that $\sigma^{n},\sigma\in\mathcal{M}_{R,M}(V)$ and the measures $\sigma^{n}$ converge VV–weakly to the measure $\sigma$. Then for every measure $\Pi\in S_{R}(\sigma)$ of the form
@@
where $(x,t)\mapsto u(x,t)$ is a Borel function from $\mathbb{R}^{d}\times[0,T]$ to $U$, there exists a sequence of measures $\Pi^{n}\in S_{R}(\sigma^{n})$ such that for every $n$ the projection of $\Pi^{n}$ on $(u,t)$ is equal to the projection of $\Pi$ on $(u,t)$ and the measures $\Pi^{n}$ converge weakly to the measure $\Pi$.
@@
$\|A(y,t,\sigma)\|+|b(y,t,\sigma)|+h^{*}(\|Q(y,t,\sigma)\|)\leq C_{2}(\sigma)V(y).$
@@
Let us verify that the measures $\Pi^{n}$ converge weakly to the measure $\Pi$. Since $\Pi^{n},\Pi\in P_{R}$ and $P_{R}$ is a compact set, it suffices to prove that if the measures $\Pi^{n}$ converge, then the limit is equal to $\Pi$. To do this it suffices to find a subsequence converging to the measure $\Pi$.
@@
$\Bigl|\int_{0}^{T}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle Q(x,t,\sigma)(u(y,t)-\widetilde{u}(y,t)),\nabla_{x}\psi(x,y)\rangle(\pi_{t}^{n}+\pi)(dxdy)\,dt\Bigr|+\\ \Bigl|\int_{0}^{T}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle Q(y,t,\sigma)(u(y,t)-\widetilde{u}(y,t)),\nabla_{y}\psi(x,y)\rangle(\pi_{t}^{n}+\pi)(dxdy)\,dt\Bigr|\leq\\ \leq 4C(\psi)\sup_{x\in B,t\in[0,T],\theta\in\mathcal{M}_{R,M}(V)}\|Q(x,t,\theta)\|\int_{0}^{T}\int_{\mathbb{R}^{d}}|u(y,t)-\widetilde{u}(y,t)|\mu_{t}(dy)\,dt.$
@@
Letting $\delta\to 0+$, we conclude that the equality $x=y$ holds for $\pi_{t}$ – almost all $(x,y)$. The idea of this proof, in particular, the choice of $H$, is well known (see, for example, [45]).
@@
Let $0<\varepsilon<\min\{2^{-1},T-T^{\prime}\}$ and $\phi$ be the standard Gaussian density on $\mathbb{R}^{d}$. We also use the function $\omega_{\varepsilon}(y)=\varepsilon^{-d}\omega(y/\varepsilon)$, where $\omega\in C_{0}^{\infty}(\mathbb{R}^{d})$, $\omega\geq 0$, $\|\omega\|_{L^{1}}=1$ and the support of $\omega$ is in $\{y\colon\,|y|<1\}$. Recall that we are given the measures $\mu_{t}$ such that
$\partial_{t}\mu_{t}=L^{*}_{\sigma,u(y,t)}\mu_{t},\quad\mu_{0}=\nu,$
where
$L_{\sigma,u(y,t)}\psi(y)={\rm trace}\bigl(A(y,t,\sigma)D^{2}\psi(y)\bigr)+\langle b(y,t,\sigma)+Q(y,t,\sigma)u(y,t),\nabla\psi(y)\rangle.$
Set
$\mu_{t}^{\varepsilon}(y):=\varepsilon\phi(y)+\frac{1-\varepsilon}{\varepsilon}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$a^{ij}_{\varepsilon}(y,t):=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}a^{ij}(z,s,\sigma)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$b^{i}_{\varepsilon}(y,t):=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}b^{i}(z,s,\sigma)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$(Qu)^{i}_{\varepsilon}(y,t):=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}Q(z,s,\sigma)u(s,z)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
$u_{\varepsilon}(y,t):=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}u(z,s)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds,$
@@
where the operator $L_{\sigma,u(z,s)}$ is applied to the function $z\mapsto\omega_{\varepsilon}(y-z)$. Note that
@@
are integrable over $\mathbb{R}^{d}\times[0,T^{\prime}]$ with respect to the measure $\mu_{t}^{\varepsilon}(y)\,dy\,dt$ and the $L^{1}$–norms of these functions are majorized by a constant independent on $\varepsilon$. Let us prove this observation. Consider the function $(Qu)^{i}_{\varepsilon}$. We have
@@
$\mathbb{B}_{n}^{\varepsilon,j}(x,y,t)=\left(\begin{array}[]{l}\zeta_{j}(x)b(x,t,\sigma^{n})+\zeta_{j}(x)Q(x,t,\sigma^{n})u_{\varepsilon}(y,t)\\ b_{\varepsilon}(y,t,\sigma)+(Qu)_{\varepsilon}(y,t,\sigma)-y\displaystyle\frac{\varepsilon\phi(y)}{\mu_{t}^{\varepsilon}(y)}\\ \end{array}\right).$
@@
over all probability measures $\eta$ with projections $\nu$ and $\nu^{\varepsilon}$ on the factors. It is known (see, for instance, [14]) that the minimum of this functional tends to zero as $\varepsilon\to 0$ because the measures $\pi_{0}^{\varepsilon}$ converge weakly to the measure $\nu\circ(x,x)^{-1}$.
According to Remark 3.4, there exists a measure $\pi^{n,\varepsilon,j}=\pi_{t}^{n,\varepsilon,j}\,dt$ such that this measure is given by a family of sub–probability measures $\pi_{t}^{n,\varepsilon,j}$ and $\pi^{n,\varepsilon,j}$ is a solution to the Cauchy problem with the operator $\mathbb{L}_{n}^{\varepsilon,j}$ and the initial condition $\pi_{0}^{\varepsilon}$.
@@
Let us consider the function $\zeta_{N}$ as above and arbitrary function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$. Substitute the function $\zeta_{N}(x)\psi(y)$ into the integral identity which defines the solution $\pi^{n,\varepsilon,j}$. Letting $N\to\infty$, we obtain that the projection of the measure $\pi^{n,\varepsilon,j}$ on $y$ is a solution to the Cauchy problem for the Fokker–Planck–Kolmogorov equation with the operator $L_{\varepsilon}$ and the initial condition $\nu^{\varepsilon}$. Since the coefficients are integrable with respect to the measure $\mu^{\varepsilon}_{t}(y)\,dy$ and the matrix $A_{\varepsilon}$ is locally Lipschitzian in $y$ and locally non-degenerate, the class of sub–probability solutions contains precisely one element $\mu^{\varepsilon}=\mu_{t}^{\varepsilon}(y)\,dy\,dt$ (see Remark 3.9). It follows that the projection of $\pi_{t}^{n,\varepsilon,j}$ on $y$ is equal to the measure $\mu_{t}^{\varepsilon}(y)\,dy$. In particular, $\pi_{t}^{n,\varepsilon,j}$ is a probability measure.
For every $t\in[0,T^{\prime}]$ let denote by $\mu^{n,\varepsilon,j}_{t}$ the projection of the measure $\pi_{t}^{n,\varepsilon,j}$ on $x$. Substitute the function $V(x)\zeta_{N}(x)\zeta_{N}(y)$ into the integral identity which defines the solution $\pi^{n,\varepsilon,j}$. Note that the coefficients depending on $x$ vanish outside a ball of sufficiently large radius. Moreover, the coefficients depending on $y$ are integrable with respect to the measure $\pi^{n,\varepsilon,j}$ since the projection of $\pi^{n,\varepsilon,j}$ on $(y,t)$ equals the measure $\mu_{t}^{\varepsilon}(y)\,dydt$. Letting $N\to\infty$, we obtain the inequality
@@
$\zeta_{j}(x)L_{\sigma^{n},u_{\varepsilon}(y,s)}V(x,s)\leq C_{L}V(x)+C_{L}Re^{Ms}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}+h(|u_{\varepsilon}(y,s)|),$
where $\beta_{V,W}$ is defined in Remark 3.2. We need the following version of Jensen’s inequality. Assume that $\Phi$ is a convex nonnegative function on $0,+\infty)$ and $\Phi(0)=0$. Let $\xi\geq 0$ be a measurable function on a measurable space $(X,\mathcal{X})$ with a sub–probability measure $\omega$. Then
@@
Let us prove this estimate. If $\omega(X)=0$, then the inequality trivially holds since $\Phi(0)=0$. Let $\omega(X)>0$. The convexity of $\Phi$ and the equality $\Phi(0)=0$ imply the inequality $\Phi(\lambda v)\leq\lambda\Phi(v)$ for every $v\geq 0$ and $0<\lambda<1$. Thus we obtain
@@
Taking into account the last estimate and the fact that the projection of $\pi^{n,\varepsilon,j}_{t}$ on $y$ does not depend on $j$, we obtain that for every $t$ the sequence $\pi^{n,\varepsilon,j}_{t}$ is tight. Hence for every $t$ one can extract a convergent subsequence $\pi^{n,\varepsilon,j_{m}}_{t}$. Arguing as in Lemma 3.3, we find a subsequence $j_{m}$ such that for every $t$ the measures $\pi^{n,\varepsilon,j_{m}}_{t}$ converge weakly to a probability measure $\pi^{n,\varepsilon}_{t}$. Moreover, the mapping $t\mapsto\pi^{n,\varepsilon}_{t}$ is continuous with respect to the weak topology. For every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ there exists a number $m_{0}$ such that for all $m>m_{0}$ and for every $(x,y,t)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\times[0,T^{\prime}]$ one has the equality